### v1

# Hypergraph Characterization of Fusion Rings

Abstract.
We present a correspondence between multiplicity-free, self-dual, fusion rings and a digraph, hypergraph pair $(D,H)$. This correspondence is used to provide a complete characterization of all fusion rings corresponding to graphical properties of $D$. Further, we exploit this correspondence to provide a complete list of all non-isomorphic, self-dual, multiplicity-free fusion rings of rank at most 8.
1. Introduction
Fusion categories arise in a number of settings in both mathematics and physics. In physics, they are part of the algebraic data necessary to describe a topological quantum field theory [22]; in computer science, they aid in understanding the properties of certain quantum computers [12, 17]; in mathematics, they axiomatize representation theory [9, 8]. Mathematically, fusion categories naturally arise from the study of representations of finite groups, quantum groups, and certain Hopf algebras [9, 8]. Fusion categories are central to many conjectures found in the literature such as the Property-F Conjecture [17], Gauging Conjecture [2], Weakly Group-Theoretical Conjecture [10], and Kaplansky’s Sixth Conjecture [13].
Fusion categories are often stratified by their rank, the number of isomorphism classes of simple objects. However, even for fixed rank fusion categories are difficult to fully characterize. At the time of writing a full classification is only known through rank 3, though partial results have been obtained up through rank 5 [18, 20, 6, 7, 14]. Part of the underlying classification problem is a lack of structure and enumeration tools. One partial solution is to characterize the Grothendieck rings of fusion categories rather than the categories themselves. This alleviates certain burdens associated with categorification and associativity, however the problem still remains difficult. In fact, there is currently no known algorithm to enumerate all candidate rings. In this paper we take a first step toward resolving this problem. To accomplish this, we will develop a correspondence between fusion rings and certain digraph/hypergraph pairs. This will enable us to apply tools from graph theory and combinatorics, fields accustomed to handling mathematical constructs with little structure. Our techniques will allow us to completely characterize certain multiplicity-free self-dual fusion rings and are extendable to cases with more complex duality structures and rings with multiplicity. Our main result is:
Theorem.
The multiplicity-free, self-dual braided fusion categories generated by an undirected, triangle-free graph $G$ are the Grothendieck equivalent to one of the following:
  1. (1)
    $\operatorname{Fib}$
  2. (2)
    $PSU(3)_{2}$,
  3. (3)
    $PSU(2)_{6}$, and
  4. (4)
    $Rep(G)$ where $G$ is an elementary abelian 2-group of order $2^{k}$.
The organization of the paper is as follows. In the next section, we introduce the notation that will be used throughout and recall the necessary definitions for algebraic topology and structural graph theory. Then in Section 3 we introduce a correspondence between fusion rings and certain digraph/hypergraph pairs. From there we begin to classify the subclasses of graphs which, by this correspondence, yield fusion rings. Our main contribution is presented in Section 4 where we provide a classification of braided fusion categories up to Grothendieck equivalence arising from triangle free undirected graphs. In Section 5 we illustrate how the correspondence developed in this work can be used to facilitate the complete enumeration of fusion rings, in particular, we give a complete listing of all multiplicity-free, self-dual fusion rings up to rank 8.
2. Preliminaries
2.1. Fusion Rings
Although our focus in this paper is the classification and enumeration of fusion rings, this study is motivated by the connections of fusion categories to both physics and mathematics. Thus, we briefly outline here the definitions associated with fusion categories and refer the interested reader to the excellent monographs [9, 8] and references therein.
A fusion category over $\mathbb{C}$, $\mathcal{C}$, is a $\mathbb{C}$-linear semisimple rigid monoidal category with finitely many simple objects and finite dimensional spaces of morphisms [18]. The fusion category is said to be braided if the monoidal product, $\otimes$, is commutative up to isomorphism with that isomorphism subject to certain consistency conditions [3]. We will denote the set of isomorphism classes of simple objects in $\mathcal{C}$ by $\operatorname{Irr}\left(\mathcal{C}\right)$ and the isomorphism classes of simple objects themselves by $X_{a}$. These simple objects will always be indexed from 00 and ordered such that $X_{0}=\mathbbm{1}$ is the monoidal unit. The dual of an object, $X$, is denoted by $X^{*}$ and the category is said to be self-dual if $X_{a}=X_{a}^{*}$ for all $X_{a}\in\operatorname{Irr}\left(\mathcal{C}\right)$. Duality induces an involution on $\operatorname{Irr}\left(\mathcal{C}\right)$ as well as the index set of $\operatorname{Irr}\left(\mathcal{C}\right)$, specifically $a^{*}$ is defined by the relationship $X_{a^{*}}=\left(X_{a}\right)^{*}$. The number of isomorphism classes of simple objects is referred to as the rank of $\mathcal{C}$. The dimensions of the hom-spaces define the fusion matrices, $N_{a}$, by $\left(N_{a}\right)_{b,c}=N_{a,b}^{c}=\dim_{\mathbb{C}}\operatorname{Hom}_{\mathcal{C}}\left(X_{a}\otimes X_{b},X_{c}\right)$. The fusion matrices are normal non-negative matrices and hence are subject to the Frobenius-Perron Theorem. The Frobenius-Perron eigenvalue of $N_{a}$ is called the Frobenius-Perron dimension or FP-dimension of $X_{a}$ and is denoted by $\operatorname{FPdim}\left(X_{a}\right)$ or more succinctly as $d_{a}$. The FP-dimension of $\mathcal{C}$ is defined by $\operatorname{FPdim}\left(\mathcal{C}\right)=\sum_{a=0}^{r-1}d_{a}^{2}$, where $r$ is the rank of the fusion category $\mathcal{C}$.
A good deal of information about a fusion category is captured by its Grothendieck ring, that is the ring formed by the equivalence classes of the objects of the category together with a certain class of relations defined by short-exact sequences in the category. Such rings can be studied in the absence of their categorical origins [19]. To this end, recall from [19], a $\mathbb{Z}_{+}$-basis of an algebra, which is free as a module over $\mathbb{Z}$, is a basis $B=\left\{X_{i}\right\}$ such that $X_{i}X_{j}=\sum_{k}N_{i,j}^{k}X_{k}$ for some $N_{i,j}^{k}\in\mathbb{N}$. In other words, the product of basis elements can be expressed as a the linear combination of the elements of the basis. The elements of the basis are called simple, the size of the basis is called the rank, and the coefficients describing the multiplication (${N_{ij}^{k}}$) are called the structure coefficients. A $\mathbb{Z}_{+}$-ring is a unital algebra over $\mathbb{Z}$ with a fixed $\mathbb{Z}_{+}$-basis. Finally, a unital based ring is a $\mathbb{Z}_{+}$-ring, $A$, with a unit $\mathbbm{1}$, a basis $X=\left\{X_{i}\right\}_{i\in I}$ with unit $\mathbbm{1}=X_{0}$, such that there exists an involution $\cdot^{*}:X\to X$ (as well as on $I$, as above) and a group homomorphism $\tau:A\to\mathbb{Z}$ which satisfy that:
  • (i)
    The induced map $\sum_{i\in I}a_{i}X_{i}\mapsto\sum_{i\in I}a_{i}X^{*}_{i}$ with $a_{i}\in\mathbb{Z}$ is an anti-involution and
  • (ii)
    For all $i,j\in I$, $\tau\left(X_{i}X_{j}\right)=\delta_{i,j^{*}}$.
The Grothendieck ring of every (braided) fusion category yields a (commutative) unital based ring. For this reason we will often refer to unital based rings as fusion rings.
Remark 1.
Henceforth we will only be interested in commutative fusion rings and so we will drop the word commutative and refer only to fusion rings.
A series of direct calculation reveals that the structure coefficients of a fusion ring are highly symmetric. For example, we note that
$\tau((X_{i}X_{j})X_{k^{*}})=\tau(\sum_{\ell}N_{ij}^{\ell}X_{\ell}X_{k^{*}})=\sum_{\ell}N_{ij}^{\ell}\delta_{\ell k}=N_{ij}^{k}$
and
$\tau((X_{i}X_{j})X_{k^{*}})=\tau(X_{i}(X_{j}X_{k^{*}}))=\tau(\sum_{\ell}N_{jk^{*}}^{\ell}X_{i}X_{\ell})=\sum_{\ell}N_{jk^{*}}^{\ell}\tau(X_{i}X_{\ell})=N_{jk^{*}}^{i^{*}}.$
Thus $N_{ij}^{k}=N_{jk^{*}}^{i^{*}}$, which is often referred to as rigidity of the fusion ring. As a consequence, we have that $N_{ij}^{0}=N_{j0}^{i^{*}}=\delta_{i^{*}j}$. Furthermore, since the duality operation is an anti-involution,
$\sum_{k}N_{ij}^{k}X_{k^{*}}=\left(\sum_{k}N_{ij}^{k}X_{k}\right)^{*}=\left(X_{i}X_{j}\right)^{*}=X_{j^{*}}X_{i^{*}}=\sum_{k^{*}}N_{j^{*}i^{*}}^{k^{*}}X_{k^{*}},$
and in particular $N_{ij}^{k}=N_{j^{*}i^{*}}^{k^{*}}$ for all $i$, $j$, and $k$. Combining these symmetries, we have that
$N_{ij}^{k}=N_{jk^{*}}^{i^{*}}=N_{k^{*}i}^{j^{*}}=N_{i^{*}k}^{j}=N_{kj^{*}}^{i}=N_{j^{*}i^{*}}^{k^{*}}.$
In fact, together with the commutativity relationship $N_{ij}^{k}=N_{ji}^{k}$, these symmetries are necessary and sufficient for the $N_{ij}^{k}$ to yield a fusion ring. This allows us to dispense with the homomorphism $\tau$, and consider fusion rings as $\mathbb{Z}$-algebras with presentation
$\mathcal{R}:=\left\langle X_{0}=\mathbbm{1},X_{1},\dots,X_{r-1}\mid X_{i}X_{j}=\sum_{k=0}^{r-1}N_{ij}^{k}X_{k},\hskip 2.84526pt\mathbbm{1}X_{i}=X_{i}\right\rangle$
where there is an involution * with $N_{ii^{*}}^{0}=1$ and the structure coefficients satisfy the three symmetry properties (rigidity, self-involution, and commutativity). In fact, not only do the structure coefficients completely determine the fusion ring, they provide a faithful representation of it. In particular, letting $a$ and $b$ be arbitrary indices, we have that
$e_{a}^{T}N_{i}N_{j}e_{b}$ $=\sum_{\ell}N_{ia}^{\ell}N_{j\ell}^{b}=\sum_{\ell}N_{ia}^{\ell}N_{b^{*}j}^{\ell*}=\sum_{s,t}N_{ia}^{s}N_{b^{*}j}^{t}\delta_{st^{*}}=\sum_{s,t}N_{ia}^{s}N_{b^{*}j}^{t}\tau(X_{s}X_{t})$
$=\tau\left(\sum_{s,t}N_{ia}^{s}N_{b^{*}j}^{t}X_{s}X_{t}\right)=\tau\left(\left(\sum_{s}N_{ia}^{s}X_{s}\right)\left(\sum_{t}N_{b^{*}j}^{t}X_{t}\right)\right)=\tau(X_{i}X_{a}X_{b^{*}}X_{j})$
and by the commutativity of the fusion ring, this is equal to
$\tau(X_{i}X_{j}X_{a}X_{b^{*}})$ $=\tau\left(\left(\sum_{s}N_{ij}^{s}X_{s}\right)\left(\sum_{t}N_{ab^{*}}^{t}X_{t}\right)\right)=\tau\left(\sum_{s,t}N_{ij}^{s}N_{ab^{*}}^{t}X_{s}X_{t}\right)=\sum_{s,t}N_{ij}^{s}N_{ab^{*}}^{t}\tau(X_{s}X_{t})$
$=\sum_{k}N_{ij}^{k}N_{ab^{*}}^{k^{*}}=\sum_{k}N_{ij}^{k}N_{ka}^{b}=\sum_{k}N_{ij}^{k}e_{a}^{T}N_{k}e_{b}.$
As a consequence, $N_{j}N_{i}=N_{i}N_{j}=\sum_{k=0}^{r-1}N_{i,j}^{k}N_{k}$ and the fusion matrices provide a faithful representation of the fusion ring.
2.2. Graphs and Hypergraphs
While our primary motivation is to study fusion rings, we will do so through the lens specifically graph theory. For this reason we will take a short detour and define the essential terminology for graphs and hypergraphs. Once this is accomplished we will establish a correspondence between fusion rings and graph/hypergraph pairs.
A (simple) graph $G$ is a ordered pair $(V,E)$ where $V$ is a finite set of vertices and $E$ is a subset of the 2-elements subsets of $V$ referred to as the edges. We say two vertices are adjacent, denoted $u\sim v$, if $\set{u,v}\in E$. A graph is said to be trivial if it contains exactly one vertex and empty if the edge set is empty. For a vertex $v$, we define its neighborhood as $N(v)=\set{u\mid\set{u,v}\in E}$. The cardinality of the neighborhood of $v$ is the degree of $v$, $\deg(v)$. We will denote by $\delta$ and $\Delta$, the minimum and maximum degree in a graph.
A subgraph of $G$ is formed by selecting a subset of vertices and edges of the graph. That is, $H=(V^{\prime},E^{\prime})$ is a subgraph of $G$ if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. A vertex-induced subgraph is formed by taking a subset of vertices $W$ and precisely those edges of $G$ with both endpoints in $W$. The subgraph of $G$ induced by $W$ is denoted $G[W]$. A path $P$ in $G$ is a sequence of vertices $(v_{0},v_{1},\dots,v_{k})$ such that one of $(v_{i},v_{i+1})$ and $(v_{i+1},v_{i})$ is an edge for all $i=0,\dots,k-1$. We define the length of a path $P=(v_{0},v_{1},\dots,v_{k})$ to be $k$. We call a path simple if it does not contain any repeated vertices. The distance between two vertices $u$ and $v$, denoted $d(u,v)$, is defined to be the length of the shortest path from $u$ to $v$. If no such path exists, the distance is defined to be infinite. The maximum distance over all pairs of vertices is called the diameter of a graph. A connected component (or component for short) of $G$ is a maximal subgraph $H$ in which any two vertices are connected by a path. We say that $G$ is connected if it contains precisely one component. We will say that a vertex is isolated if its connected component has precisely one vertex.
Similarly to a graph, a directed graph or digraph, is an ordered pair $(V,E)$ where $E$ is a subset of $V\times V$. Much of the terminology above extends naturally to digraphs with additional consideration taken for the direction of the edges. For example, instead of the neighborhood of a vertex $v$ we consider the in- and out-neighborhoods of a vertex, $N^{-}(v)=\set{u\mid(u,v)\in E}$ and $N^{+}(v)=\set{u\mid(v,u)\in E}$, respectively. The in- and out-degrees of a vertex are defined analogously and denoted $\deg^{-}(v)$ and $\deg^{+}(v)$. An edge of the form $(v,v)$ is called a loop and we will say a vertex is trivial if it is an isolated vertex without a self-loop.
A hypergraph is a generalization of an undirected graph where the edges can join, not just pairs, but any number of vertices. Formally, a hypergraph is a pair $H=(V,E)$ where $V$ is a set of vertices and $E\subset\mathcal{P}(V)$ is a set of nonempty subsets of $V$. The elements of $E$ are called hyperedges. In the special case where all the hyperedges have size $k$, we say that $H$ is $k$-uniform. In particular, a simple graph is a $2$-uniform hypergraph. For further graph theoretic terms we refer the reader to [4].
3. Fusion Ring/Digraph Correspondence
In this section we will establish a correspondence between a certain class of fusion rings which are multiplicity-free (i.e., $N_{ij}^{k}\leq 1$ for all $i,j,k$) and self-dual (i.e., the involution associated with the fusion ring is the identity map) and certain digraph/hypergraph pairs. This correspondence will enable the majority of our study.
We note that the restriction to considering self-dual fusion rings induces additional symmetries on the structure coefficients, in particular,
$N_{ij}^{k}=N_{ji}^{k}=N_{ik}^{j}=N_{ki}^{j}=N_{jk}^{i}=N_{kj}^{i}.$
Thus, it is easy to see that for fixed rank $r$ all multiplicity-free, self-dual fusion rings can be determined by considering all $\binom{r}{3}+2\binom{r}{2}+r$ unique possible entries in the fusion matrices. However, this approach has several implicit inefficiencies that significantly limit its effectiveness. Specifically, it fails to account for the dependencies between the entries resulting from the commutation of the fusion matrices and it fails to account for the $\sim(r-1)!$ isomorphic copies of each assignment resulting from the relabeling of the elements of the basis. By rephrasing the problem of guessing entries of the fusion matrices as guessing a digraph and a hypergraph we can partially address both these issues. This is accomplished by exploiting the commutation relationship and exploiting extant tools for dealing with graph isomorphism [16].
More concretely, we note that there are $2\binom{r}{2}+r$ possible edges in a $r$-vertex digraph with self-loops and $\binom{r}{3}$ possible edges in a 3-uniform hypergraph on $r$ vertices. Thus we can build a representation of all potential fusion matrices by specifying the edges of a directed graph, $D$, with self-loops on $\set{0,1,\ldots,r-1}$ and the edges of a 3-uniform hypergraph, $H$, on the same set. Specifically, the digraph and hypergraph encode the fusion rules as follows:
  • for $i\in\set{0,1,\ldots,r-1}$, there is a self-loop at $i$ in $D$ if and only if $N_{ii}^{i}=1$,
  • for $\set{i,j}\subseteq\set{0,1,\ldots,r-1}$ there is a directed edge $(i,j)$ in $D$ if and only if $N_{ii}^{j}=N_{ij}^{i}=N_{ji}^{i}=1$, and
  • for $\set{i,j,k}\subseteq\set{0,1,\ldots,r-1}$ there is a hyperedge $\set{i,j,k}$ in $H$ if and only if $N_{ij}^{k}=N_{ji}^{k}=N_{ik}^{j}=N_{ki}^{j}=N_{jk}^{i}=N_{kj}^{i}=1$.
We note that the entries of $N_{i}$ complete describe the edges incident with the node $i$, in particular, there is a self-loop at $i$ if and only if $\left(N_{i}\right)_{ii}=1$, a directed edge $(i,j)$ if and only if $\left(N_{i}\right)_{jj}=1$, a directed edge $(j,i)$ if and only if $\left(N_{i}\right)_{ij}=\left(N_{i}\right)_{ji}=1$, and a hyperedge $\set{i,j,k}$ if and only if $\left(N_{i}\right)_{jk}=\left(N_{i}\right)_{kj}=1.$ In particular, the edges incident with 00 are completely determined by the convention that $X_{0}=\mathbbm{1}$ and thus $N_{0}=I$. Specifically, the only edges incident to 00 are a self-loop and a directed edge from 00 to $i\in[r-1]=\set{1,\ldots,r-1}.$ Thus for the remainder, we will focus on the correspondence between a self-dual, multiplicity-free rank $r$ fusion ring and a digraph/hypergraph pair on $[r-1]$. In order to illustrate the correspondence between fusion rule, fusion matrix, a graphical presentations of a fusion ring, in Figure 1, we present all three presentations for the fusion ring $\operatorname{Rep}\left(\mathbb{Z}_{3}^{2}\rtimes\mathbb{Z}_{2}\right)$.
For an arbitrary digraph, hypergraph pair $(D,H)$ we will say that $(D,H)$ generates a fusion ring if the associated matrices $\set{N_{i}}$ are the fusion matrices for some fusion ring. We will further say that a digraph $D$ generates a fusion ring if there exists some hypergraph $H$ such that the pair $(D,H)$ generates a fusion ring. Finally, we say that $D$ (resp. $(D,H)$) generates a fusion category if $D$ (resp. $(D,H)$) generates the Grothendieck ring of the category.
$\otimes$ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$
$X_{1}$ $\mathbbm{1}+X_{1}+X_{5}$ $X_{3}+X_{4}$ $X_{2}+X_{4}$ $X_{2}+X_{3}$ $X_{1}$
$X_{2}$ $X_{3}+X_{4}$ $\mathbbm{1}+X_{2}+X_{5}$ $X_{1}+X_{4}$ $X_{1}+X_{3}$ $X_{2}$
$X_{3}$ $X_{2}+X_{4}$ $X_{1}+X_{4}$ $\mathbbm{1}+X_{3}+X_{5}$ $X_{1}+X_{2}$ $X_{3}$
$X_{4}$ $X_{2}+X_{3}$ $X_{1}+X_{3}$ $X_{1}+X_{2}$ $\mathbbm{1}+X_{4}+X_{5}$ $X_{4}$
$X_{5}$ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $\mathbbm{1}$
(a) Fusion Rule Presentation
$\left[\begin{matrix}0&0&0&1&0&0\\ 0&0&1&0&1&0\\ 0&1&0&0&1&0\\ 1&0&0&1&0&1\\ 0&1&1&0&0&0\\ 0&0&0&1&0&0\end{matrix}\right]$$X_{3}$$\left[\begin{matrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{matrix}\right]$$\mathbbm{1}$$\left[\begin{matrix}0&0&0&0&1&0\\ 0&0&1&1&0&0\\ 0&1&0&1&0&0\\ 0&1&1&0&0&0\\ 1&0&0&0&1&1\\ 0&0&0&0&1&0\end{matrix}\right]$$X_{4}$$\left[\begin{matrix}0&1&0&0&0&0\\ 1&1&0&0&0&1\\ 0&0&0&1&1&0\\ 0&0&1&0&1&0\\ 0&0&1&1&0&0\\ 0&1&0&0&0&0\end{matrix}\right]$$X_{1}$$\left[\begin{matrix}0&0&0&0&0&1\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 1&0&0&0&0&0\end{matrix}\right]$$X_{5}$$\left[\begin{matrix}0&0&1&0&0&0\\ 0&0&0&1&1&0\\ 1&0&1&0&0&1\\ 0&1&0&0&1&0\\ 0&1&0&1&0&0\\ 0&0&1&0&0&0\end{matrix}\right]$$X_{2}$
(b) Fusion Matrix Presentation
[…]
(c) Graphical Presentation
Figure 1. Presentations of the Fusion Ring $\operatorname{Rep}\left(\mathbb{Z}_{3}^{2}\rtimes\mathbb{Z}_{2}\right)$.
It is worth noting that this digraph/hypergraph construction can easily be extended to non-multiplicity-free fusion rings by allowing for positive integer weights on the edges and hyperedges. However, this is beyond of the scope of this work, and we will not consider it further.
4. Fusion Rings Generated by Undirected Graphs
In this section we will make use of the correspondence established above. Specifically we will consider natural sets of digraph, e.g., undirected graphs, connected graphs, and triangle free graphs, and ask when such graphs can generate a fusion ring. Each graph result will be paired with the implication on the side of fusion categories. Many of the fusion category statements are verbose due to notions in graph theory not being naturally expressible in the language of fusion categories, e.g., neighborhood. Nonetheless, our results provide a new perspective on the classification and enumeration of fusion categories. The power of this approach is displayed in our main result, Theorem 3, where we are able to provide a complete characterization of fusion categories (up to Grothendieck equivalence) generated by triangle free undirected graphs.
Throughout this section we will assume that the digraph $D$ is undirected, that is if $(i,j)$ is an edge in $D$, then so is $(j,i)$. In order to emphasize this assumption we will use $G$ to denote the digraph and simply refer to a graph, hypergraph pair. In terms of the fusion ring this condition is equivalent to:
$N_{ii}^{j}=N_{ji}^{i}=N_{ij}^{i}=N_{jj}^{i}=N_{ji}^{j}=N_{ij}^{j}\quad\text{for all $i,j\in[r-1]$.}$
For notational brevity we introduce the following for a given pair $\left(G,H\right)$:
[…]
$w_{ij}$ $\equiv\textrm{number of $k$ such that $\set{i,j,k}$ is a hyperedge.}$.
Using the fact that the fusion matrices commute (in particular, that $\left(N_{i}N_{j}\right)_{ij}=\left(N_{j}N_{i}\right)_{ij}$) it is easy to see that the $w_{ij}$’s and the $e_{ij}$’s are related via
$w_{ij}=1+\sum_{k\in[r-1]}e_{ik}e_{jk}-2e_{ij}.$ (1)
Combinatorially, this equation can be interpreted as saying that if $i$ and $j$ are adjacent, then the number of hyperedges is one less than the number of triangles containing $\set{i,j}$ in $G$ plus the number of loops on $i$ or $j$. If $i$ and $j$ are not adjacent, then the number of hyperedges containing $\set{i,j}$ is one more than the number of common neighbors of $i$ and $j$ in $G$.
4.1. Empty Graphs
We first consider the situation where the graph $G$ has no edges, that is, $G$ consists of some number of isolated vertices. In this case, we can impose additionally structure on the associated hypergraph $H$, namely, that it is a $3$-uniform hypergraph where every pair of vertices is in precisely one hyperedge. This is more commonly referred to as a Steiner triple system [15].
Theorem 1.
Let $G$ be the empty graph on $n$ vertices and let $H$ be such that $(G,H)$ generates a fusion ring, then $n=2^{k}-1$ for some $k$ and $H$ is a Steiner triple system of order $12^{k}-$ 1. Furthermore, the $H$ that generates the fusion ring is unique up to isomorphism.
Proof.
Since $G$ is an empty graph, for any $1\leq i<j\leq n$ we have that $N_{ii}^{j}=N_{jj}^{i}=0$. As a consequence, for all such $i$ and $j$, $(N_{i}N_{j})_{ij}=\sum_{k}N_{ii}^{k}N_{jk}^{j}=N_{ii}^{0}N_{j0}^{0}=1$. Observing that by the commutativity we have that that $1=(N_{j}N_{i})_{ij}=\sum_{k}N_{ji}^{k}N_{ik}^{j}$ and that this counts the number of hyperedges containing $\set{i,j}$ plus the number of directed edges between $i$ and $j$, we conclude that there is a unique hyperedge containing $\set{i,j}$. That is, $H=([n],E)$ is a Steiner triple system. As Steiner triple systems are known to exist if and only if $n\equiv 1,2\pmod{6}$, (see for instance [15]) this significantly restricts the possible self-dual, multiplicity-free fusion rings that can be generated by the empty graph. Indeed, letting $H$ be an arbitrary Steiner triple system generates a set of valid fusion coefficients which respects rigidity and that the duality operation is an anti-involution. However, not all such Steiner triple systems will resulting in a commuting family of fusion matrices.
In order to understand which Steiner triple systems result in a commuting family of fusion matrices, we define for each $i\in[n]$ the graph $G_{i}$ to be the graph whose adjacency matrix is $N_{i}$. We note the fact that $N_{ia}^{b}=N_{ib}^{a}$ implies that this is $G_{i}$ is a graph rather than a digraph. Observing that $(N_{i})_{i0}=1$ and that for all $i,j\in[n]$ there is a unique $k$ such that $(N_{i})_{jk}=1$, we conclude that each $N_{i}$ has precisely one non-zero entry in each row, and hence $G_{i}$ is a perfect matching. Finally, noting that for each (not necessarily distinct) $j,k\in\set{0,\ldots,n}$ there is precisely one $i\in\set{0,\ldots,n}$ such that $(N_{i})_{jk}=1$ and thus $\sum_{i=0}^{r}N_{i}=J$ where $J$ is the $(n+1)\times(n+1)$ matrix of ones. Now, as $N_{0}=I$ we have that $\sum_{i=1}^{r}N_{i}=J-I$ which is the adjacency matrix of the complete graph on $n+1$ vertices. Thus we have that the $G_{i}$ form a partition of the edges of $K_{n+1}$ into $n$ commuting perfect matchings. By [1, Theorem 1], such a decomposition exists if and only if $n+1=2^{k}$ for some $k\in\mathbb{N}$. Further, the pair $(G,H)$ generates a fusion ring if and only if $H$ is Steiner triple system defined from a commuting collection of perfect matchings.
In order to prove uniqueness of $H$ (up to relabeling) we show that in any decomposition of $K_{2^{k}}$ into commuting perfect matchings there exists a collection of $k$ of these perfect matchings whose union forms a hypercube of dimension $k$. Furthermore, for every perfect matching outside this set, the edge set is uniquely specified by any edge in the perfect matching.
By way of induction suppose that $M_{1},\ldots,M_{t}$ is a collection of commuting perfect matchings on $[2^{k}]$ whose union consists of $2^{k-t}$ disjoint $t$-dimensional hypercubes. Now let $M$ be a perfect matching that commutes with and is disjoint from all of the $M_{i}$. From [1], this implies that $M\cup M_{i}$ is a collection of disjoint 4-cycles for all $i$. Let $S$ be the vertex set of one of the $t$-dimensional hypercubes in $\cup_{i}M_{i}$. We claim that either $M[S]$ has zero edges or $2^{t-1}$ edges. To this end suppose $e$ is an edge in $M[S]$ and without loss of generality label the vertices of $S$ with length $t$ binary strings such that one endpoint of $e$ is the binary 00-string string and for all $i$ the endpoints of the edges in $M_{i}$ differ only in the $i^{\textrm{th}}$ component. Now we note that since $M$ and $M_{i}$ commute, if $e$ is an edge in $M[S]$, then the edge with $i^{\textrm{th}}$ component of both endpoints of $e$ changed is also in $M[S]$. We will abuse notation and denote this edge by $e+e_{i}$. By a similar argument we have that for $j\neq i$, $(e+e_{i})+e_{j}$ is also in $M[S]$. Noting that this operation is commutative, the action of each matching $M_{i}$ is of order two, and that there is a collection of actions which reverse the labels of the endpoints of $e$, we have that $M[S]$ contains at least $2^{t-1}$ edges. In fact, it contains exactly this many as $M$ is a perfect matching and $\left|S\right|=2^{t}$.
Thus, for every hypercube in $\cup M_{i}$, there are at most $12^{t}-t-$ 1 commuting perfect matchings on $[2^{k}]$ containing an edge internal to the hypercube. As there are $2^{k-t}$ such hypercubes, there are at most $2^{k-t}\left(2^{t}-t-1\right)=2^{k}-(t+1)2^{k-t}$ commuting hypercubes which have an edge in any of the hypercubes in $\cup M_{i}$. Since there are at $12^{k}-t-$ 1 unused matchings, there is some matching $M_{t+1}$ which does not contain an edge internal to any hypercube in $\cup M_{i}$. By a similar argument as above, $M_{t+1}\cup\bigcup_{i}M_{i}$ forms $2^{k-t-1}$ hypercubes of dimension $t+1$ on $[2^{k}]$. Repeating this process yields $M_{1},\ldots,M_{k}$, a collection of $k$ commuting perfect matchings on $[2^{k}]$ whose union is a hypercube. The above argument also gives that every matching outside $\set{M_{1},\ldots,M_{t}}$ is completely determined by a single edge, as desired. ∎
Algebraic Corollary 1.
If $\mathcal{C}$ is a multiplicity-free, self-dual, braided fusion category generated by the empty graph on $n$ vertices, then $n=2^{k}-1$ for some $k\in\mathbb{Z}$ and $\mathcal{C}$ is Grothendieck equivalent to $\operatorname{Rep}\left(\mathbb{Z}_{2}^{k}\right)$.
4.2. Simple Components
In order to further our understanding of which undirected graphs $G$ can generate a fusion ring we will next consider the non-trivial components $G$ which do not contain a self-loop.
Lemma 1.
Let $G$ be an undirected graph that generates a fusion ring and let $C$ be a nontrivial component of $G$ which contains no self-loops. For every vertex $v\in C$, the minimum degree of $C[N(v)]$ is at least 2.
Proof.
Suppose not. Then there exists a vertex $v$ which has a neighbor $u$ with degree at most 1 in $C[N(v)]$. In fact, $u$ must have a exactly one neighbor in $C[N(v)]$ since otherwise, by Equation 1, $\{u,v\}$ would participate in $-1$ hyperedges. Let $x$ be the unique vertex that is adjacent to both $u$ and $v$. Then noting that $w_{uv}=0$ by Equation 1, we have that
$(N_{v}N_{u})_{xv}$ $=\sum_{k}N_{vx}^{k}N_{uk}^{v}=N_{vx}^{u}e_{uv}+e_{vx}e_{vu}=e_{vx}e_{vu}=1,\textrm{ while }$
$(N_{v}N_{u})_{vx}$ $=\sum_{\ell}N_{vv}^{\ell}N_{u\ell}^{x}\geq e_{vu}e_{ux}+e_{vx}e_{xu}=2$
a contradiction. ∎
Algebraic Corollary 2.
If $\mathcal{C}$ is a multiplicity-free, self-dual, braided fusion category such that $N_{i,j}^{j}=N_{j,i}^{i}$ for all $i,j\neq 0$ and $N_{i,i}^{i}=0$, then for all non-invertible $X\in\operatorname{Irr}\left(\mathcal{C}\right)$, $X\otimes X$ contains at least three non-isomorphic subobjects.
It is an easy consequence of Lemma 1 that every vertex and edge in such a loopless component participates in at least 2 triangles. Consequently the minimum degree of such a component is at least 3. In fact, with care we can show that the minimum degree of such a component is at least 4.
Lemma 2.
Suppose that $G$ is an undirected graph which generates a fusion ring and let $C$ be a non-trivial component of $G$ with no self-loops, then the minimum degree of $C$ is at least 4.
Proof.
Let $H$ be a hypergraph such that $(G,H)$ generates a fusion ring and let $C$ be a non-trivial simple component of $G$ with minimum degree at most $3$. As a consequence of Lemma 1, the component $C$ contains a vertex $\ell$ of degree exactly 3. Furthermore, $C[N(\ell)\cup\{\ell\}]$ is a complete graph on 4 vertices, $\{i,j,k,\ell\}.$ By Equation 1 we have that $w_{\ell i}=w_{\ell j}=w_{\ell k}=1$. Suppose now that $\set{i,j,\ell}$ is the unique hyperedge involving $i$ and $\ell$, and note that
$(N_{\ell}N_{i})_{ij}=\sum_{u}N_{\ell i}^{u}N_{iu}^{j}=e_{\ell i}N_{i\ell}^{j}+e_{i\ell}e_{ij}+N_{\ell i}^{j}e_{ji}=3,$
while
$(N_{\ell}N_{i})_{ji}=\sum_{v}N_{\ell j}^{v}N_{iv}^{i}=e_{j\ell}e_{ij}+e_{\ell j}e_{i\ell}+N_{\ell j}^{i}e_{ii}=2,$
which is a contradiction to the commutativity of $N_{\ell}$ and $N_{i}$. As $i$ and $j$ were arbitrary, we have that $\set{i,k,\ell}$ and $\set{j,k,\ell}$ are also not hyperedges in $H$.
Thus, there is some vertex $s\notin\set{i,j,k,\ell}$ such that $\set{s,i,\ell}$ is a hyperedge. Now note that
$(N_{\ell}N_{j})_{\ell i}=\sum_{u}N_{\ell\ell}^{u}N_{ju}^{i}=e_{\ell i}e_{ij}+e_{\ell j}e_{ji}+e_{\ell k}N_{jk}^{i}=2+N_{jk}^{i}$
and
$(N_{\ell}N_{j})_{i\ell}=\sum_{v}N_{\ell i}^{v}N_{jv}^{\ell}=e_{\ell i}e_{\ell j}+e_{i\ell}N_{ji}^{\ell}+N_{\ell i}^{s}N_{js}^{\ell}=1+N_{j\ell}^{s}.$
Thus, by the symmetry of $N_{\ell}N_{i}$ we have that $\{i,j,k\}\notin E(H)$ and $\{s,j,\ell\}\in E(H)$. A similar argument shows that $\{s,k,\ell\}\in E(H)$. Figure 2 illustrates the required edges and hyperedges thus far.
$\ell$$i$$j$$k$$s$
$\ell$$i$$j$$k$$s$
Figure 2. Initial edges and hyperedges required by vertex of degree 3.
As s≁not-$s\not\sim\ell$ and $w_{\ell s}\geq 3$, we have by Equation 1 that $\ell$ and $s$ have at least two common neighbors in $G$. Without loss of generality suppose that $i$ and $j$ are common neighbors of $\ell$ and $s$. At this point, we have that either $s\not\sim k$ and $w_{\ell s}=3$ or $s\sim k$ and there exists a vertex $t\not\in\{i,j,k\}$ such that $\{\ell,s,t\}\in E(H)$. Now considering
$(N_{\ell}N_{k})_{ks}=\sum_{u}N_{\ell k}^{u}N_{ku}^{s}=e_{\ell k}N_{k\ell}^{s}+e_{k\ell}e_{ks}+N_{\ell k}^{s}e_{sk}=1+2e_{ks},$
and
$(N_{\ell}N_{k})_{sk}=\sum_{v}N_{\ell s}^{v}N_{kv}^{k}=N_{\ell s}^{i}e_{ki}+N_{\ell s}^{j}e_{kj}+N_{\ell s}^{t}e_{kt}+N_{\ell s}^{k}e_{kk}=2+N_{\ell s}^{t}e_{kt},$
we have that $1+e_{ks}=2+N_{\ell s}^{t}e_{k,t}$. Thus, $s\sim k$, $\{\ell,s,t\}\in E(H)$, and $k\sim t$. But as we now know that $k\sim s$, this implies that $i$ and $j$ were chosen arbitrarily. Repeating the previous argument for $\set{i,k}$ and $\set{j,k}$ and using that $w_{\ell s}=4$, we get that $i\sim t$ and $j\sim t$. As a result we have the edges and hyperedges depicted in Figure 3.
$\ell$$i$$j$$k$$s$$t$
$\ell$$i$$j$$k$$s$$t$
Figure 3. Edges and hyperedges required by neighbors of vertex of degree 3.
Now consider $i$ and $j$. By Equation 1, $w_{ij}\geq 3$. Further, from the above, we know that $\{i,j,k\}$ and $\{i,j,\ell\}$ are not hyperedges. Thus, there exists a vertex $a\notin\{\ell,k,s,t\}$ in the graph such that $\{i,j,a\}$ is a hyperedge. Now note that
$(N_{\ell}N_{i})_{\ell a}=\sum_{u}N_{\ell\ell}^{u}N_{iu}^{a}\geq e_{\ell j}N_{ij}^{a}\geq 1,$
and
$(N_{\ell}N_{i})_{a\ell}=\sum_{v}N_{\ell a}^{v}N_{iv}^{\ell}=e_{\ell a}e_{\ell i}+N_{\ell a}^{i}e_{iv}+N_{\ell a}^{s}N_{is}^{\ell}=0,$
contradicting the symmetry of $N_{\ell}N_{i}$. Thus, we may conclude that the minimum degree in $C$ is at least 4. ∎
Corollary 1.
If $\mathcal{C}$ is a self-dual, multiplicity-free, braided fusion category then either there exists a non-trivial simple object $X$ such that $\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)\ncong 0$, or for all non-invertible $Y\in\operatorname{Irr}\mathcal{C}$ $Y\otimes Y$ contains at least 4 non-isomorphic non-trivial subobjects.
Remark 1.
In [11], they show that an undirected graph with minimum degree at least 4 contains either $K_{5}$ or $K_{2,2,2}$ as a minor. Further, by Kuratowski’s Theorem, a graph with a $K_{5}$ minor is nonplanar. Thus a graph coming from a fusion ring which contains an undirected, simple, nontrivial component is either nonplanar or contains a $K_{2,2,2}$ minor.
4.3. Non-simple Components
In order to understand the behavior of the non-simple components of an undirected graph generating a fusion ring it will be helfpul to rephrase implications of Equation 1 in terms of the local loop structure of graph.
Lemma 3.
Let $G$ be an undirected graph that generates a fusion ring.
  1. (1)
    If $i\not\sim j$ in $G$, then $w_{ij}=1+|N(i)\cap N(j)|.$
  2. (2)
    If $i\sim j$ in $G$ and $(N(i)\cap N(j))\setminus\{i,j\}=\varnothing$, then
    • $w_{ij}=0$ and there is a loop in $G$ on precisely one of $i$ and $j$, or
    • $w_{ij}=1$ and there is a loop on both vertices in $G$.
  3. (3)
    If $i\sim j$ in $G$ and both vertices are loopless, then they must share a common neighbor.
Our first observation shows that if an undirected graph $G$ generates a fusion ring then it has at most one component containing loops.
Lemma 4.
Let $G$ be an undirected graph that generates a fusion ring, then for any two looped vertices $i$ and $j$ in $G$, $d(i,j)\leq 2$.
Proof.
By way of contradiction, suppose that $i$ and $j$ are looped vertices in $G$ and $d(i,j)\geq 3$. By Lemma 3, $w_{ij}=1$. Now let $\{i,j,k\}$ be the unique hyperedge containing $i$ and $j$ and note that
$(N_{i}N_{j})_{ik}=\sum_{u}N_{ii}^{u}N_{ju}^{k}\geq e_{ii}N_{ji}^{k}=1,$
and
$(N_{i}N_{j})_{ki}=\sum_{v}N_{ik}^{v}N_{jv}^{i}=e_{ik}e_{ij}+e_{ki}N_{jk}^{i}+N_{ik}^{j}e_{ji}=e_{ki}$
Thus, by the symmetry of $N_{i}N_{j}$ we must have the edge $\{i,k\}$ in $G$. Further, by comparing $(N_{i}N_{j})_{jk}$ and $(N_{i}N_{j})_{kj}$, $\{j,k\}$ must also be present, contradicting that $d(i,j)\geq 3$. Thus either $i\sim j$ or $d(i,j)=2$, as desired. ∎
Algebraic Corollary 3.
If $\mathcal{C}$ is a multiplcity-free, self-dual, braided fusion category, then for $X\ncong Y\in\operatorname{Irr}\left(\mathcal{C}\right)$ either $Y$ is a subobject of $X\otimes X$, or $X\otimes X$ and $Y\otimes Y$ have a common simple subobject.
As a consequence of this result, there is only one non-simple component in graph $G$ which generates a fusion ring, and furthermore that component has diameter at most 2. It might seem that the diameter 2 constraint would significantly limit the number of simple graphs that need be considered as the base for the non-simple component in $G$, however a classic result of Erdős and Renyí gives that a vanishingly small fraction of graphs have diameter greater than 2, see [5].
Lemma 5.
Let $G$ be an undirected graph that generates a fusion ring. If $G$ contains an induced path, $i\sim j\sim k$ where $\set{i,j}$ is not in a triangle, then $i$ and $j$ both have self loops.
Proof.
By Lemma 3, it suffices to show that there is a hyperedge containing $i$ and $j$. To that end we note that
$1=N_{ii}^{j}N_{kj}^{j}\leq\sum_{u}N_{ii}^{u}N_{ku}^{j}=(N_{i}N_{k})_{ij}=(N_{i}N_{k})_{ji}=\sum_{v}N_{ij}^{v}N_{kv}^{i}\leq e_{ji}N_{kj}^{i}+\sum_{v\neq i,j}N_{ij}^{v}=N_{kj}^{i}+w_{ij}.$
Thus there is a hyperedge containing both $i$ and $j$, as desired. ∎
Algebraic Corollary 4.
Let $\mathcal{C}$ be a multiplicity-free, self-dual, braided fusion category. If $X,Y,Z$ are pairwise non-isomorphic simple objects such that $Y$ is a subobject of $X\otimes X$ and $Z$ is a subobject of $Y\otimes Y$, but $X\otimes X$ and $Y\otimes Y$ have no common subobjects other than potentially $X$ and $Y$, then $\dim\operatorname{Hom}_{\mathcal{C}}\left(X\otimes X,X\right)=\dim\operatorname{Hom}_{\mathcal{C}}\left(Y\otimes Y,Y\right)=1$.
Note that this implies that any triangle-free component with at least 3 vertices of a graph generating a fusion ring has a self-loop at every vertex.
4.4. Fusion Rings Generated by Triangle-Free Graphs
We now turn to the principle result of this section, a complete characterization of the undirected, triangle-free graphs which generate a fusion ring.
Theorem 2.
The undirected, triangle-free graphs which generate fusion rings are:
  1. (1)
    a single vertex with a self loop,
  2. (2)
    a single edge with one self loop,
  3. (3)
    a single edge with self loops on each vertex and an isolated loopless vertex, and
  4. (4)
    a collection of $12^{k}-$ 1 isolated loopless vertices.
Proof.
Let $G$ be a triangle-free, undirected graph which generates a fusion ring. In the case that $G$ is the empty graph, by Theorem 1, $G$ has order $12^{k}-$ 1 and has a unique hypergraph $H$ associated with which it generates a fusion ring. Thus we may assume that $G$ contains at least one edge and let $T$ be the maximal collection of nontrivial triangle-free components. By Lemma 3 every edge in $T$ has a self-loop on at least one of the vertices. But then by Lemma 4 the diameter of $T$ is at most $4$ and in particular $T$ is a single component. In particular, $G$ consists of $T$ and some number of trivial (that is, isolated without a self-loop) vertices.
Suppose now that $T$ contains at least 3 vertices. Since $T$ is connected and contains at least $3$ vertices, for any edge $\set{i,j}$ in $T$, there exists some vertex $k$ such that either $i\sim k$ or $j\sim k$. Since $T$ is triangle-free, by Lemma 5 both $i$ and $j$ have a self-loop. As $\set{i,j}$ is an arbitrary edge, this implies that all vertices in $T$ have a self-loop and by Lemma 4 $T$ has diameter at most 2. In fact, since $T$ is triangle-free and has at least three vertices, this implies that the diameter of $T$ is exactly 2. Furthermore, by Lemma 3 for every edge $\set{i,j}\in E(T)$, $w_{ij}=1$. In particular, any hypergraph which, with $G$, generates a fusion ring is non-empty. Let $H$ be one such hypergraph and let $f=\{i,j,k\}$ be an arbitrary edge in $H$. Since $T$ is triangle-free, there are at most two edges of $T$ in $G[f]$. Suppose then that there are precisely two edges from $T$ in $f$, namely $i\sim j$ and $j\sim k$. Since $i,j,k$ are all in $T$, we have that $e_{ii}=e_{jj}=e_{kk}=1$ and, by Lemma 3, $w_{ik}=w_{kj}=1$. Furthermore, the hyperedge indicated by $w_{ij}$ and $w_{jk}$ is $f$. But now we have
$(N_{i}N_{j})_{ik}=\sum_{u}e_{iu}N_{ju}^{k}=e_{ii}N_{ji}^{k}+e_{ik}e_{kj}+e_{ij}e_{jk}=2.$
and
$(N_{i}N_{j})_{ki}=\sum_{v}N_{ik}^{v}N_{jv}^{i}=e_{ki}N_{jk}^{i}+e_{ik}e_{ij}+N_{ik}^{j}e_{ji}=1,$
which contradicts the symmetry of $N_{i}N_{j}$. Thus, for any hyperedge $f$, $G[f]$ must contain at most one edge from $T$.
We note that by Lemma 3 and the non-trivial nature of $T$, there exists some hyperedge $f$ such that $G[f]$ contains one edge of $T$. More specifically, suppose that $i$ and $j$ are adjacent vertices in $T$, then $w_{ij}=1$. Let $k$ be the third vertex in the unique hyperedge containing $i$ and $j$. Suppose $k$ lies in $T$. Since $T$ is triangle free and has diameter 2, there must exist a vertex $x$ such that $i,j,x,k$ forms an induced a path of length 4 in $T$. Now since $w_{xj}=1$ by Lemma 3, we have that
$(N_{x}N_{i})_{xj}=\sum_{u}e_{xu}N_{iu}^{j}=e_{xj}e_{ji}+e_{xk}N_{ik}^{j}+e_{xi}e_{ij}=2,$
and
$(N_{x}N_{i})_{jx}=\sum_{v}N_{xj}^{v}N_{iv}^{x}\leq e_{xj}e_{xi}+e_{jx}N_{ij}^{x}+\sum_{s\neq j,x}N_{xj}^{s}N_{is}^{x}\leq 1,$
again contradicting the symmetry of $N_{x}N_{i}$. Thus for any hyperedge containing an edge in $T$, the third vertex in the hyperedge must not belong to $T$.
Now suppose that $i$ and $j$ are two non-adjacent vertices in $T$. Since $T$ is triangle free and has diameter 2, there is a vertex $k$ such that $i\sim k$ and $j\sim k$. But then we have that,
$1=e_{ik}e_{jk}\leq\sum_{u}N_{ii}^{u}N_{ju}^{k}=(N_{i}N_{j})_{ik}=(N_{i}N_{j})_{ki}=\sum_{v}N_{ik}^{v}N_{jv}^{i}=\sum_{v\neq 0,i,j,k}N_{ik}^{v}N_{jv}^{i}.$
Thus there is some vertex $\ell$ such that both $\{i,k,\ell\}$ and $\{i,j,\ell\}$ are hyperedges in $H$. But since $i\sim k$, we know that $\ell$ is not in $T$, and in particular, $i$ and $\ell$ have no common neighbors. Thus, by Lemma 3, $w_{i\ell}=1$, a contradiction. Hence there are no non-adjacent vertices in $T$, and so $T$ has at most two vertices. That is, $T$ is a looped vertex or an edge with either one or two self-loops.
First suppose that $T$ is a looped vertex $x$. By Lemma 3 any trivial vertices must occur in exactly one hyperedge with $x$. This would require two isolated vertices, $i$ and $j$, being contained in a hyperedge with a looped vertex. However, this implies that $01=(N_{i}N_{j})_{ix}=(N_{i}N_{j})_{xi}=$ 0, a contradiction. Thus, if $T$ is a single looped vertex, then $T$ is the entire graph $G$. Next, suppose that $T$ is an edge $\{x,y\}$ with at least one loop. Without loss of generality, first suppose that only $x$ is looped. Since $w_{xy}=0$, additional trivial vertices must occur, in pairs, in exactly one hyperedge with $x$. Again we have that $T$ is the entire graph $G$. Finally suppose that both $x$ and $y$ are looped. By Lemma 3, $w_{xy}=1$ and thus at least one trivial vertex must be present. However, if more than one trivial vertex is present, a hyperedge of the form $\{x,i,j\}$ must also be present, where $i$ and $j$ are trivial vertices. Thus, such an edge must occur with exactly one trivial vertices. ∎
From the fusion rules corresponding to the graphs in Theorem 2 it is easy verify the following theorem.
Theorem 3.
The multiplicity-free, self-dual braided fusion categories generated by an undirected, triangle-free graph $G$ are the Grothendieck equivalent to one of the following:
  1. (1)
    $\operatorname{Fib}$
  2. (2)
    $PSU(3)_{2}$,
  3. (3)
    $PSU(2)_{6}$, and
  4. (4)
    $Rep(G)$ where $G$ is an elementary abelian 2-group of order $2^{k}$.
5. Low Rank Fusion Rings
One of the advantages of the digraph representation of fusion rings is that it allows for the use of already existing tools to facilitate the complete enumeration. For example, in this section we use the utilities distributed McKay and Piperno’s graph isomorphism tool nauty/Traces [16] to generate all non-isomorphic digraphs on a given number of vertices. Each digraph then yields restrictions on the associated hypergraph through the relationship between $e_{ij}$ and $w_{ij}$. From simple combinatorial considerations on the hypergraph, a significant fraction of digraphs can be shown not to generate a fusion ring. For the remaining digraphs, a direct search through the associated hypergraphs will yield all digraph-hypergraph pairs which generate a fusion ring. A complete table of those pairs through rank 8 is provided below. In addition to the fusion rings, we provide the Grothendieck class for the fusion rings, if known. To facilitate this procedure we make the following observation:
Observation 1.
If two digraphs, $D_{1}$ and $D_{2}$, generate fusion rings $F_{1}$ and $F_{2}$, there is a digraph $D$ that is a super graph of $D_{1}\times D_{2}$ which generates $F_{1}\boxtimes F_{2}$. More concretely let $D_{1}^{\prime}$ be formed from $D_{1}$ by adding the vertex 00 which has a self-loop and directed edges from every vertex in $D_{1}$ to 00. Then $D=D_{1}^{\prime}\times D_{2}^{\prime}-\set{(0,0)}$ and $D$ generates the fusion ring $F_{1}\boxtimes F_{2}$.
Remark 2.
It is perhaps interesting to note that in the tables below, the tuple
$\left(\sum_{i}N_{ii}^{i},\sum_{i}\sum_{j\neq i}N_{ii}^{j},\sum_{\left|\set{i,j,k}\right|=3}N_{ij}^{k}\right)$
uniquely determines the fusion ring. Furthermore, in examining the rank 5 fusion rings we see that the more algebraically natural tuples $\left(\sum_{i}\operatorname{Tr}\!\left(N_{i}\right),\sum_{\left|\set{i,j,k}\right|=3}N_{ij}^{k}\right)$ or $\left(\sum_{i}\operatorname{Tr}\!\left(N_{i}\right),\sum_{i,j,k}N_{i,j}^{k}\right)$ fail to distinguish between all fusion rings. We are not so bold as to conjecture that this holds in general, but it may be a fruitful direction of future research.
Table 1. Rank 2 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$\operatorname{Sem}$[21]
$1$ $\operatorname{Fib}$[21]
Table 2. Rank 3 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$(1,2)$ $\operatorname{Ising}$[20]
$1$ $(1,2)$ $\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[20]
$1$ $(1,2)$, $(2,1)$ $\operatorname{PSU}(3)_{2}$
Table 3. Rank 4 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$(1,2,3)$ $\operatorname{Sem}^{2}$[21]
$(1,2)$, $(1,3)$, $(2,1)$, $(2,3)$ $\operatorname{PSO}(5)_{6}$[6]
$2$ $(1,2)$ $(1,2,3)$ $\operatorname{Sem}\boxtimes\operatorname{Fib}$[21]
$1$, $2$ $(1,2)$, $(2,1)$ $(1,2,3)$ $\operatorname{PSU}(2)_{6}$[6]
$1$, $2$ $(1,2)$, $(1,3)$, $(2,1)$, $(3,2)$ $(1,2,3)$ $\operatorname{PSU}(2)_{5}$[6]
$1$, $2$, $3$ $(1,2)$, $(1,3)$ $(1,2,3)$ $\operatorname{Fib}^{2}$ [21]
Table 4. Rank 5 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$(1,2)$, $(1,3)$, $(1,4)$ $(2,3,4)$ $\mathrm{TY}(\nicefrac{{\mathbb{Z}}}{{2\mathbb{Z}}}\times\nicefrac{{\mathbb{Z}}}{{2\mathbb{Z}}},\chi,\nu)$
$(1,3)$, $(1,4)$, $(2,1)$, $(2,4)$, $(3,2)$, $(3,4)$ $(1,2,3)$ $\mathcal{C}(\operatorname{\mathfrak{so}}_{7},e^{\nicefrac{{\pi i}}{{14}}},14)_{ad}$
$1$ $(1,2)$, $(1,3)$, $(1,4)$ $(2,3,4)$ not categorifiable[9]
$1$ $(1,4)$, $(2,1)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$ $\mathcal{C}(\operatorname{\mathfrak{so}}_{3},e^{\nicefrac{{\pi i}}{{6}}},6)$
$1$, $3$ $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$
$1$, $3$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$ $(1,2,3)$, $(2,3,4)$
$2$, $3$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(3,2)$ $(1,2,3)$, $(2,3,4)$
$1$, $2$, $3$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(2,3,4)$
$1$, $2$, $3$ $(1,4)$, $(2,1)$, $(2,3)$, $(3,1)$, $(3,2)$ $(1,2,3)$, $(2,3,4)$
$1$, $2$, $3$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,4)$, $(3,2)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
Table 5. Rank 6 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$(1,5)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{Ising}$ [21]
$(1,2)$, $(1,5)$, $(2,1)$, $(2,5)$, $(3,1)$, $(3,2)$, $(4,1)$, $(4,2)$ $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
$2$ $(1,2)$, $(1,5)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[21, 20]
$4$ $(1,3)$, $(1,4)$, $(1,5)$, $(2,5)$, $(3,4)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{Ising}$[21]
$2$ $(1,2)$, $(1,4)$, $(2,4)$, $(3,2)$, $(4,2)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(3,4,5)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}\left(3\right)_{2}$[21]
$1$ $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
$2$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,3)$, $(2,4)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$
$1$ $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$
$1$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$, $(5,2)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$
$1$, $2$ $(1,2)$, $(1,5)$, $(2,1)$, $(2,5)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$
$1$, $2$ $(1,2)$, $(1,4)$, $(2,1)$, $(2,4)$, $(3,1)$, $(3,2)$, $(4,1)$, $(4,2)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
$1$, $2$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$
$1$, $2$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,5)$, $(3,4)$ $(1,2,3)$, $(1,2,4)$, $(3,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{Rep}\left(\mathfrak{S}_{3}\right)$[20]
$1$, $3$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,5)$, $(5,3)$ $(1,2,3)$, $(1,2,5)$, $(1,3,4)$, $(2,4,5)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(3)_{2}$ [20]
$1$, $2$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,2)$, $(3,4)$, $(4,2)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(3,4,5)$
$1$, $2$, $4$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(3,1)$, $(3,2)$, $(3,4)$, $(4,2)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(2,3,5)$, $(3,4,5)$
$1$, $2$, $3$, $4$ $(1,5)$, $(2,5)$, $(3,5)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$
$1$, $2$, $3$, $4$ $(1,5)$, $(2,4)$, $(2,5)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$
$1$, $2$, $3$, $4$ $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,4)$, $(4,1)$, $(4,2)$, $(4,3)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(2,3,4)$, $(3,4,5)$
$1$, $2$, $4$, $5$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,5)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(2,4,5)$
Table 6. Rank 7 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,6)$, $(3,1)$, $(3,6)$ $(1,2,3)$, $(2,3,4)$, $(2,3,5)$, $(4,5,6)$
$(1,5)$, $(1,6)$, $(2,1)$, $(2,6)$, $(3,2)$, $(3,6)$, $(4,3)$, $(4,6)$, $(5,4)$, $(5,6)$ $(1,2,4)$, $(1,3,4)$, $(1,3,5)$, $(2,3,5)$, $(2,4,5)$
$(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,5)$, $(3,6)$, $(4,3)$, $(4,6)$, $(5,4)$, $(5,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(3,4,5)$
$3$ $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,6)$, $(3,1)$, $(3,6)$ $(1,2,3)$, $(2,3,4)$, $(2,3,5)$, $(4,5,6)$
$1$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,6)$, $(3,1)$, $(3,6)$ $(1,2,3)$, $(2,3,4)$, $(2,3,5)$, $(4,5,6)$
$3$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,6)$, $(3,1)$, $(3,6)$ $(1,2,3)$, $(2,3,4)$, $(2,3,5)$, $(4,5,6)$
$2$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,5)$, $(3,6)$, $(4,3)$, $(4,6)$, $(5,4)$, $(5,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(3,4,5)$
$2$, $3$ $(1,2)$, $(1,3)$, $(1,6)$, $(2,3)$, $(3,2)$, $(4,3)$, $(5,3)$ $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(2,3,6)$, $(2,4,5)$, $(4,5,6)$
$2$, $3$ $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,6)$, $(3,1)$, $(3,2)$, $(3,6)$ $(1,2,3)$, $(2,3,4)$, $(2,3,5)$, $(4,5,6)$
$1$, $5$ $(1,6)$, $(2,1)$, $(2,4)$, $(2,5)$, $(3,1)$, $(3,4)$, $(3,5)$, $(4,1)$, $(4,3)$, $(4,5)$, $(5,1)$, $(5,3)$ $(1,2,3)$, $(1,4,5)$, $(2,3,4)$, $(2,3,5)$, $(2,3,6)$, $(2,4,5)$, $(4,5,6)$
$1$, $2$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,5)$, $(3,6)$, $(4,3)$, $(4,6)$, $(5,4)$, $(5,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(3,4,5)$
$3$, $4$, $5$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(2,3)$, $(2,5)$, $(3,5)$, $(4,5)$, $(5,4)$ $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(2,3,4)$, $(2,3,6)$, $(4,5,6)$
$1$, $2$, $3$ $(1,6)$, $(2,1)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,1)$, $(4,3)$, $(5,1)$, $(5,3)$ $(1,2,3)$, $(1,4,5)$, $(2,3,4)$, $(2,3,5)$, $(2,3,6)$, $(2,4,5)$, $(4,5,6)$
$1$, $2$, $3$, $5$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,5)$, $(4,3)$, $(4,5)$, $(5,3)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(2,3,4)$, $(2,3,6)$, $(2,4,5)$, $(4,5,6)$
$1$, $2$, $3$, $5$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,5)$, $(4,1)$, $(4,3)$, $(4,5)$, $(5,3)$, $(6,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,3,4)$, $(1,3,5)$, $(2,3,4)$, $(2,3,6)$, $(2,4,5)$, $(3,4,6)$, $(4,5,6)$
$1$, $2$, $3$, $5$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,2)$, $(4,3)$, $(4,5)$, $(4,6)$, $(5,4)$, $(5,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(2,3,4)$, $(2,3,5)$, $(2,3,6)$
$1$, $2$, $3$, $4$, $5$ $(1,4)$, $(1,5)$, $(1,6)$, $(2,1)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,1)$, $(3,2)$, $(3,4)$, $(3,5)$, $(4,3)$, $(4,5)$, $(5,3)$, $(5,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,3,4)$, $(1,3,5)$, $(2,3,4)$, $(2,3,5)$, $(2,3,6)$, $(2,4,5)$, $(4,5,6)$
$1$, $2$, $3$, $4$, $5$, $6$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(2,3)$, $(2,4)$, $(2,6)$, $(3,2)$, $(3,5)$, $(3,6)$, $(4,2)$, $(4,3)$, $(4,5)$, $(5,2)$, $(5,4)$, $(5,6)$, $(6,3)$, $(6,4)$, $(6,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,3,4)$, $(1,3,5)$, $(1,3,6)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(2,3,5)$, $(2,4,6)$, $(2,5,6)$, $(3,4,5)$, $(3,4,6)$
Table 7. Rank 8 Fusion Rings
Loops Arcs Hyperedges Grothendick Class
$(1,2,3)$, $(1,4,5)$, $(1,6,7)$, $(2,4,6)$, $(2,5,7)$, $(3,4,7)$, $(3,5,6)$ $\operatorname{Sem}^{3}$[21]
$(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(5,1)$, $(5,7)$, $(6,1)$, $(6,7)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(2,5,6)$, $(3,4,7)$, $(3,5,6)$, $(4,5,6)$ $\operatorname{Sem}\boxtimes\operatorname{PSO}(5)_{6}$ [21, 6]
$(1,4)$, $(1,7)$, $(2,6)$, $(2,7)$, $(3,5)$, $(3,7)$, $(4,2)$, $(4,7)$, $(5,1)$, $(5,7)$, $(6,3)$, $(6,7)$ $(1,2,3)$, $(1,2,6)$, $(1,3,4)$, $(1,5,6)$, $(2,3,5)$, $(2,4,5)$, $(3,4,6)$, $(4,5,6)$
$(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,6)$, $(2,7)$, $(4,1)$, $(4,6)$, $(4,7)$, $(5,1)$, $(5,6)$, $(5,7)$, $(6,3)$, $(6,7)$, $(7,3)$, $(7,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(2,4,5)$, $(3,4,5)$, $(4,5,6)$, $(4,5,7)$
$7$ $(1,7)$, $(2,7)$, $(3,7)$ $(1,2,3)$, $(1,2,4)$, $(1,3,5)$, $(1,6,7)$, $(2,3,6)$, $(2,5,7)$, $(3,4,7)$, $(4,5,6)$ $\operatorname{Fib}\boxtimes\operatorname{Sem}^{2}$[21]
$7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(2,6)$, $(2,7)$, $(6,3)$, $(6,7)$, $(7,3)$ $(1,2,3)$, $(1,2,4)$, $(1,2,6)$, $(1,2,7)$, $(3,4,5)$, $(4,6,7)$, $(5,6,7)$
$4$ $(2,4)$, $(3,1)$, $(3,5)$, $(5,1)$, $(5,3)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,4)$, $(2,3,7)$, $(2,5,6)$, $(3,4,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSO}(5)_{6}$[21, 6]
$5$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,6)$, $(2,7)$, $(4,3)$, $(4,6)$, $(5,3)$, $(6,3)$, $(6,7)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(4,5,6)$, $(4,5,7)$, $(5,6,7)$
$6$, $7$ $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(6,7)$, $(7,6)$ $(1,2,3)$, $(1,2,6)$, $(1,2,7)$, $(1,4,6)$, $(1,5,7)$, $(2,4,7)$, $(2,5,6)$, $(3,4,5)$, $(3,6,7)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}(2)_{6}$[21, 6]
$6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,2)$, $(5,6)$, $(6,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$
$6$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(3,7)$, $(5,7)$, $(6,5)$, $(6,7)$, $(7,6)$ $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(2,3,5)$, $(2,3,6)$, $(2,4,7)$, $(3,4,5)$, $(5,6,7)$ $\operatorname{Sem}\boxtimes\operatorname{PSU}(2)_{5}$[21, 6]
$5$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(4,1)$, $(4,5)$, $(4,6)$, $(5,1)$, $(5,6)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(2,3,6)$, $(2,4,5)$, $(3,4,5)$, $(4,5,7)$
$4$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,2)$, $(3,5)$, $(4,2)$, $(5,2)$, $(5,6)$, $(6,2)$, $(6,3)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(4,5,6)$
$5$, $6$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(3,6)$, $(7,5)$, $(7,6)$ $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,3,7)$, $(1,4,7)$, $(2,3,7)$, $(2,4,5)$, $(3,4,6)$, $(5,6,7)$ $\operatorname{Fib}^{2}\boxtimes\operatorname{Sem}$[21]
$1$, $5$, $6$ $(1,5)$, $(1,6)$, $(5,1)$, $(5,6)$, $(6,1)$, $(6,5)$, $(7,1)$, $(7,5)$, $(7,6)$ $(1,2,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,5,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,2)$, $(5,6)$, $(6,2)$, $(7,1)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$
$1$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,7)$, $(7,1)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(3,5,6)$, $(4,5,6)$, $(5,6,7)$
$1$, $4$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,2)$, $(3,5)$, $(4,2)$, $(5,2)$, $(5,6)$, $(6,2)$, $(6,3)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,7)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(4,5,6)$
$1$, $4$, $7$ $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$4$, $5$, $6$ $(1,4)$, $(1,5)$, $(2,4)$, $(2,5)$, $(4,5)$, $(4,6)$, $(5,4)$, $(5,6)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,5)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,4,6)$, $(1,5,7)$, $(1,6,7)$, $(2,4,7)$, $(2,5,6)$, $(2,6,7)$, $(3,4,5)$, $(3,6,7)$, $(4,5,7)$, $(4,6,7)$, $(5,6,7)$
$1$, $4$, $7$ $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $3$, $5$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(3,1)$, $(3,4)$, $(3,5)$, $(3,7)$, $(4,1)$, $(4,3)$, $(4,5)$, $(5,1)$, $(5,3)$, $(6,1)$, $(6,3)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,6)$, $(2,4,5)$, $(3,4,6)$, $(3,4,7)$, $(3,5,6)$, $(3,6,7)$, $(4,5,6)$, $(4,5,7)$, $(5,6,7)$
$1$, $4$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $2$, $3$, $4$ $(1,5)$, $(2,5)$, $(3,5)$, $(4,5)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,6,7)$, $(4,6,7)$, $(5,6,7)$
$1$, $5$, $6$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(5,1)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,5)$, $(7,6)$ $(1,2,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,6,7)$, $(3,5,7)$, $(4,5,6)$, $(5,6,7)$
$4$, $5$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,6)$, $(2,7)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,1)$, $(5,2)$, $(5,4)$, $(5,6)$, $(5,7)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,6)$ $(1,2,4)$, $(1,2,6)$, $(1,2,7)$, $(1,3,6)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(1,5,7)$, $(2,3,7)$, $(2,4,5)$, $(2,5,6)$, $(3,4,5)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$
$1$, $2$, $3$, $5$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,6)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(3,7)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,5)$, $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,5,6)$, $(1,5,7)$, $(2,3,5)$, $(2,6,7)$, $(3,4,5)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(5,6,7)$
$1$, $2$, $4$, $6$, $7$ $(1,4)$, $(4,1)$, $(5,2)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,6)$ $(1,2,6)$, $(1,3,4)$, $(1,5,7)$, $(1,6,7)$, $(2,3,5)$, $(2,4,7)$, $(3,6,7)$, $(4,5,6)$, $(4,6,7)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(2)_{6}$ [21, 6]
$1$, $2$, $3$, $4$, $7$ $(1,6)$, $(2,6)$, $(3,6)$, $(4,6)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,5,7)$, $(2,3,4)$, $(2,5,7)$, $(3,5,7)$, $(4,5,7)$, $(5,6,7)$
$1$, $4$, $5$, $6$, $7$ $(1,3)$, $(1,5)$, $(2,4)$, $(2,5)$, $(2,6)$, $(3,5)$, $(5,1)$, $(6,1)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,4,7)$, $(1,6,7)$, $(2,3,4)$, $(2,3,6)$, $(2,5,7)$, $(2,6,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$ $\operatorname{Fib}\boxtimes\operatorname{PSU}(2)_{5}$[21, 6]
$1$, $4$, $5$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,4)$, $(2,5)$, $(2,7)$, $(3,7)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,4)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,4)$, $(6,5)$, $(6,7)$, $(7,5)$ $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,4)$, $(1,3,5)$, $(1,4,5)$, $(1,4,6)$, $(1,5,6)$, $(1,6,7)$, $(2,3,5)$, $(2,3,7)$, $(2,4,6)$, $(2,5,6)$, $(3,4,6)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$
$1$, $2$, $3$, $4$, $5$, $7$ $(1,6)$, $(2,6)$, $(3,6)$, $(4,6)$, $(5,1)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,7)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,4)$, $(1,3,4)$, $(1,5,7)$, $(2,3,4)$, $(2,5,7)$, $(3,5,7)$, $(4,5,7)$, $(5,6,7)$
$1$, $2$, $3$, $4$, $5$, $7$ $(1,5)$, $(1,6)$, $(1,7)$, $(2,5)$, $(2,6)$, $(2,7)$, $(3,5)$, $(3,6)$, $(3,7)$, $(4,6)$, $(5,4)$, $(5,7)$, $(7,4)$, $(7,5)$ $(1,2,3)$, $(1,2,4)$, $(1,2,5)$, $(1,2,7)$, $(1,3,4)$, $(1,3,5)$, $(1,3,7)$, $(2,3,4)$, $(2,3,5)$, $(2,3,7)$, $(4,5,7)$, $(5,6,7)$
$1$, $3$, $4$, $5$, $6$, $7$ $(1,4)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,4)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,4,7)$, $(1,5,6)$, $(3,4,5)$, $(3,4,6)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $2$, $4$, $5$, $6$, $7$ $(1,4)$, $(1,5)$, $(1,6)$, $(2,4)$, $(2,5)$, $(2,6)$, $(4,1)$, $(4,2)$, $(4,5)$, $(4,6)$, $(4,7)$, $(5,1)$, $(5,4)$, $(5,6)$, $(6,2)$, $(6,4)$, $(6,5)$, $(7,1)$, $(7,2)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,4)$, $(1,2,5)$, $(1,2,6)$, $(1,2,7)$, $(1,3,5)$, $(1,4,6)$, $(1,4,7)$, $(1,5,7)$, $(1,6,7)$, $(2,3,6)$, $(2,4,5)$, $(2,4,7)$, $(2,5,7)$, $(2,6,7)$, $(3,4,7)$, $(4,5,6)$, $(4,5,7)$, $(4,6,7)$, $(5,6,7)$
$1$, $3$, $4$, $5$, $6$, $7$ $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,7)$, $(4,2)$, $(5,1)$, $(5,2)$, $(5,6)$, $(5,7)$, $(6,1)$, $(6,2)$, $(6,3)$, $(6,7)$, $(7,1)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(3,4,5)$, $(3,4,6)$, $(3,5,6)$, $(3,5,7)$, $(3,6,7)$, $(4,5,6)$, $(5,6,7)$
$1$, $2$, $3$, $4$, $5$, $6$, $7$ $(1,4)$, $(1,5)$, $(2,4)$, $(2,6)$, $(3,5)$, $(3,6)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,3)$, $(1,2,7)$, $(1,3,7)$, $(1,4,5)$, $(1,6,7)$, $(2,3,7)$, $(2,4,6)$, $(2,5,7)$, $(3,4,7)$, $(3,5,6)$ $\operatorname{Fib}^{3}$ [21]
$1$, $2$, $3$, $4$, $5$, $6$, $7$ $(1,2)$, $(1,3)$, $(1,4)$, $(1,5)$, $(2,1)$, $(2,3)$, $(2,4)$, $(2,6)$, $(3,1)$, $(3,2)$, $(3,5)$, $(3,6)$, $(4,1)$, $(4,3)$, $(4,5)$, $(4,6)$, $(5,2)$, $(5,3)$, $(5,4)$, $(5,6)$, $(6,1)$, $(6,2)$, $(6,4)$, $(6,5)$, $(7,1)$, $(7,2)$, $(7,3)$, $(7,4)$, $(7,5)$, $(7,6)$ $(1,2,4)$, $(1,2,5)$, $(1,2,7)$, $(1,3,5)$, $(1,3,6)$, $(1,3,7)$, $(1,4,6)$, $(1,4,7)$, $(1,5,6)$, $(1,5,7)$, $(1,6,7)$, $(2,3,4)$, $(2,3,6)$, $(2,3,7)$, $(2,4,5)$, $(2,4,7)$, $(2,5,6)$, $(2,5,7)$, $(2,6,7)$, $(3,4,5)$, $(3,4,6)$, $(3,4,7)$, $(3,5,7)$, $(3,6,7)$, $(4,5,7)$, $(4,6,7)$, $(5,6,7)$